OpenStudy (thesecret20111):

Can someone please explain the solution to this indefinite integral? :)

OpenStudy (thesecret20111):

I've been supplied the solution but I'm not sure how it works so an explanation would be great :)

OpenStudy (freckles):

u=3x+1 then du=3 dx or solving for dx gives 1/3 du=dx

OpenStudy (freckles):

\[\int\limits_{}^{}u^5 \frac{1}{3} du \\ \frac{1}{3} \int\limits u^5 du\]

OpenStudy (freckles):

can you evaluate that ?

OpenStudy (thesecret20111):

Ohhhh I see how that works! Thanks so much!! :D

OpenStudy (freckles):

np